### v3

Abstract.
For a matroid of rank $r$ and a non-negative integer $k$, an element is called $k$-loose if every circuit containing it has size greater than $r-k$. Zaslavsky and the author characterized all binary matroids with a $1$-loose element. In this paper, we establish a sharp linear bound on the size of a binary matroid, in terms of its rank, that contains a $k$-loose element. A matroid is called $k$-paving if all its elements are $k$-loose. Rajpal showed that for a prime power $q$, the rank of a $GF(q)$-matroid that is $k$-paving is bounded. We provide a bound on the rank of $GF(q)$-matroids that are cosimple and have two $k$-loose elements. Consequently, we strengthen the result of Rajpal by providing a bound on the rank of $GF(q)$-matroids that are $k$-paving. Additionally, we provide a bound on the size of binary matroids that are $k$-paving.
Key words and phrases:
$k$-paving matroid, $k$-loose elements, Reed-Muller codes
1991 Mathematics Subject Classification:
05B35, 94B65
1. Introduction
The notation and terminology follow [2] unless stated otherwise. All graphs and matroids considered here are finite and simple. For a matroid of rank $r$ and a non-negative integer $k$, an element $t$ of $M$ is called $k$-loose if every circuit of $M$ that contains $t$ has size greater than $r-k$. A matroid $M$ is called $k$-paving if every element of $M$ is $k$-loose. Note that a matroid $M$ is paving if and only if $M$ is $1$-paving. Acketa [1] provided a characterization of all $1$-paving matroids that are binary, while Oxley [3] determined all ternary $1$-paving matroids. Zaslavsky and the author characterized all binary matroids with a $1$-loose element [6]. In Section 2, we show that the size of a binary matroid that contains a $k$-loose element is linear in terms of its rank. In particular, we provide the following sharp bound.
Theorem 1.1.
For $r\geq 5$ and $0\leq k\leq\frac{r-2}{3}$, let $M$ be a simple binary matroid of rank $r$ with no coloops and a $k$-loose element. Then $|E(M)|\leq 2^{k}(r-k+1)$.
We say a matroid $M$ is a $GF(q)$-matroid if $M$ is representable over the field of $q$ elements. Section 3 considers $GF(q)$-matroids that have two $k$-loose elements. We show that the rank of such a matroid is bounded unless the two $k$-loose elements form a cocircuit of size two. The following is the precise statement.
Theorem 1.2.
Let $M$ be a simple $GF(q)$-matroid with no coloops that has two $k$-loose elements $e$ and $f$. Then $r(M)\leq(q+1)(k-1)+2q$, or $\{e,f\}$ is a cocircuit of $M$.
An immediate consequence of Theorem 1.2 is the following.
Corollary 1.3.
Let $M$ be a simple $GF(q)$-matroid with no coloops that is $k$-paving. Then $M$ is a circuit or $r(M)\leq(q+1)(k-1)+2q$.
Rajpal [5, Proposition 8] demonstrated the existence of a rank bound for $k$-paving matroids representable over $GF(q)$. Corollary 1.3 strengthens this result by giving an explicit rank bound for $GF(q)$-matroids that are $k$-paving. Observe that paving matroids are precisely $1$-paving matroids. Therefore by Corollary 1.3, a paving matroid representable over $GF(q)$ has rank at most $2q$, unless it is a circuit.
A binary $k$-paving matroid is maximal if it is not a proper minor of any other binary $k$-paving matroid. Rajpal [5] established a connection between maximal binary $k$-paving matroids and Reed-Muller codes, proving the following.
Proposition 1.4.
For $k\geq 1$, let $G$ be a generator matrix of the Reed-Muller code $R(1,k+2)$. Then the vector matroid of $G$ is a maximal binary $k$-paving matroid.
Additionally, Rajpal determined all binary $2$-paving matroids. In Section 4, we provide a bound on the size of binary $k$-paving matroids.
Theorem 1.5.
Let $M$ be a simple binary $k$-paving matroid of rank $r$, where $M$ is not a circuit and $r\geq k+4$, and let $t=\lfloor\frac{3k+1-r}{2}\rfloor$. Then $r(M)\leq 3k+1$, and
$|E(M)|\leq(r+1)+\sum_{i=0}^{t}\binom{k}{i}.$
Note that every simple binary matroid with rank at most $k+2$ is $k$-paving, and every simple triangle-free binary matroid of rank $k+3$ is $k$-paving. This includes binary affine matroids of rank $k+3$. To demonstrate the assistance Theorem 1.5 provides in finding all $k$-paving binary matroids, we apply it to $3$-paving binary matroids and obtain the following result.
Corollary 1.6.
Let $M$ be a simple binary $3$-paving matroid of rank at least seven. Then, $r(M)\leq 10$ and $|E(M)|\leq 13$. In particular, the size of $M$ is at most $12,13,11,$ and $12$ when the rank of $M$ is $7,8,9,$ and $10,$ respectively.
2. $k$-loose elements in binary matroids
Theorem 1.1 provides a sharp linear bound on the size of a binary matroid containing a $k$-loose element, in terms of its rank. Specifically, if a binary matroid $M$ has a $1$-loose element, then the size of $M$ is at most $2r$. If $M$ has a $2$-loose element, then the size of $M$ does not exceed $44r-$ 4.
Proof of Theorem 1.1.
Let $e$ be a $k$-loose element of $M$. Note that if $k=0$, then by [4, Lemma 3.1], $M$ is a circuit, and the result follows. Therefore $k\geq 1$. Suppose that $e$ is not $(k-1)$-loose. It follows that there is a basis $B=\{e_{1},\ldots,e_{r}\}$ of $M$ such that, in the standard binary representation $P=[I_{r}|Q]$ of $M$ with respect to $B$, the column $e^{P}$ labeled by $e$ has its first $k$ entries equal to zero, while all other entries are non-zero. We call the set of indices of the non-zero entries of a column $f^{P}$ in $Q$ the support of $f^{P}$, denoted as $\text{supp}(f^{P})$. The entries of $f^{P}$, excluding the first $k$ entries, are called the root entries of $f^{P}$, and the set of indices of the non-zero root entries of $f^{P}$ is the root support of $f^{P}$. Let $Q-e^{P}$ denote the columns in $Q$ excluding $e^{P}$.
2.0.1.
For a column $f^{P}$ in $Q-e^{P}$, the sum of the number of zeroes in the first $k$ entries of $f^{P}$ and the number of non-zero root entries of $f^{P}$ is at most $k+1$.
Suppose otherwise. Then $e^{P}+f^{P}$ has at least $k+2$ zeroes, so at most $r-(k+2)$ columns from $I_{r}$ can be added to $e^{P}+f^{P}$ to obtain a column with all entries equal to zero. It follows that $e$ is in a circuit of size at most $r-k$, a contradiction. Therefore 2.0.1 holds.
Let $F$ be a set of all columns in $Q-e^{P}$ such that, for any pair $f^{P},g^{P}$ in $F$, the sum $f^{P}+g^{P}$ has its first $k$ entries equal to zero. It is clear that the first $k$ entries of all columns in $F$ are identical.
2.0.2.
Let $f^{P},g^{P}$ be columns in $F$. Then, the symmetric difference of their supports, $supp(f^{P})\bigtriangleup supp(g^{P})$, contains at most two elements.
Suppose not. Then $f^{P}+g^{P}$ has its first $k$ entries equal to zero and contains at least three non-zero entries. It follows that $e^{P}+f^{P}+g^{P}$ has at least $k+3$ zeroes so it can be combined linearly over GF(2) with at most $r-(k+3)$ columns in $I_{r}$ to obtain the column that has all the entries zero. Therefore $e$ is in a circuit of size at most $r-k$, a contradiction. Thus 2.0.2 holds.
2.0.3.
$|F|\leq r-k+1$. Moreover, if the columns in $F$ have exactly one non-zero entry in the first $k$ entries, then $|F|\leq r-k$.
Choose a column $f^{P}$ in $F$ such that the root support of $f^{P}$, say $S$, has the largest possible size. Suppose $|S|=s$. It follows by 2.0.1 that $s\leq k+1$. In view of 2.0.2, a column $g^{P}$ in $F-f^{P}$ can fall into one of the following three categories, namely,
  1. (i)
    $\text{supp}(g^{P})\subseteq\text{supp}(f^{P})$ with $|\text{supp}(f^{P})-\text{supp}(g^{P})|=1$,
  2. (ii)
    $\text{supp}(g^{P})\subseteq\text{supp}(f^{P})$ with $|\text{supp}(f^{P})-\text{supp}(g^{P})|=2$, or
  3. (iii)
    $\text{supp}(g^{P})\not\subseteq\text{supp}(f^{P})$.
We call these columns type A, type B, and type C, respectively. By 2.0.2, it follows that the number of type B columns is at most $\text{max}\{3,s-1\}$. Note that if $s\leq 1$, there is no type B column. If $s=2$, there is at most one type B column. It follows that the number of type B columns does not exceed $s$.
First, suppose that $F$ contains a type C column. We say an entry of a column $g^{P}$ is in $S$ if the index of the entry is in $S$. Similarly, an entry of $g^{P}$ is outside $S$ if its index is not in $S$. Since the size of the root support of any column in $F$ is at most $s$, it follows from 2.0.2 that a type C column has precisely one non-zero root entry outside $S$ and exactly one zero entry in $S$. Observe that if $F$ contains two type C columns such that their corresponding entries in $S$ are not identical, then by 2.0.2, all their corresponding entries outside $S$ must be identical. In this case, the number of type C columns is at most $s$. On the other hand, if all type C columns in $F$ have identical entries in $S$, then the number of type C columns is at most $r-(s+k)$. Therefore, the number of columns of type C in $F$ is at most $\text{max}\{s,r-(s+k)\}$. Given that $s\leq k+1$ and $k\leq\frac{r-2}{3}$, it follows that the number of type C columns in $F$ is at most $r-(s+k)$. Since $F$ contains a type C column, it is immediate from 2.0.2 that there is at most one type A column in $F$. Furthermore, if a type A column exists, then we have at most $s-1$ type B columns. It follows that the sum of the number of type A columns and the number of type B columns does not exceed $s$. Therefore the size of $F$ is at most $r-(s+k)+s+1=r-k+1$.
Now, suppose that no type C column exists. By 2.0.2, it follows that if a type B column exists, then the number of type A columns is at most $2$ so $F$ has size at most $s+3$. Since $s\leq k+1,r\geq 5$, and $k\leq\frac{r-2}{3}$, it follows that $s+3\leq r-k+1$. Finally note that the number of type A columns is at most $s$ so $|F|\leq r-k+1$.
If the columns in $F$ have exactly one non-zero entry in the first $k$ entries, then by 2.0.1, $s$ is at most two. If $s=1$, then all the columns in $F-f^{P}$ are type C; if a column is of type A, then it is parallel to some element in $I_{r}$, a contradiction since $M$ is simple. Since the number of type C columns is at most $r-(s+k)$, it follows that $|F|$ is at most $r-k$. Now suppose that $s=2$. Note that all columns in $F-f^{P}$ are either type A or type C; a type B column contradicts that $M$ is simple. It is clear that the number of type A columns is at most two and the number of type C columns is $r-(2+k)$. However, if $F$ contains a type C column, then, by 2.0.2, the number of type A columns is at most one. Therefore $|F|$ is at most $\text{max}\{3,r-k\}$. Since $k\leq\frac{r-2}{3}$, it follows that $|F|\leq r-k$. Thus 2.0.3 holds.
Observe that if there is a column $h^{P}$ in $Q-e^{P}$ that has its first $k$ entries equal to zero, then by 2.0.1, the column $h^{P}$ has at most one non-zero root entry so $h^{P}$ has at most one non-zero entry. Since $M$ is simple, this is a contradiction. It follows that there are at most $12^{k}-$ 1 sets $F$ in $Q-e^{P}$, and $k$ of these have exactly one non-zero entry in the first $k$ entries. Therefore, by 2.0.3, $|E(M)|\leq k(r-k)+(2^{k}-1-k)(r-k+1)+(r+1)=2^{k}(r-k+1)$.
If $e$ is $(k-1)$-loose, we choose the smallest non-negative integer $t$ such that $e$ is $t$-loose. By the same argument we obtain $|E(M)|\leq 2^{t}(r-t+1)$. Since $t\leq k-1,r\geq 5$, and $k\leq\frac{r-2}{3}$, it follows that $2^{t}(r-t+1)\leq 2^{k}(r-k+1)$. ∎
The following proposition demonstrates that the bound in Theorem 1.1 is sharp.
Proposition 2.1.
For $r\geq 2$ and $20\leq k\leq r-$ 2, there exists a simple binary matroid of rank $r$ with a $k$-loose element, whose size is $2^{k}(r-k+1)$.
Proof.
Let $M$ be a binary matroid represented by the matrix $P=[I_{r}|Q]$ over $GF(2)$ such that $Q$ contains a column $e$ whose first $k$ entries are zero while the remaining entries are non-zero. We now describe the other columns of $Q$, demonstrate that $M$ has size $2^{k}(r-k+1)$, and show that the element corresponding to the column $e$ is $k$-loose in $M$.
For a column $f$ in $Q$, we refer to the column obtained by removing the first $k$ entries of $f$ as the root vector of $f$. The columns in $Q-e$ can be partitioned into disjoint sets $F$, where every column in $F$ has identical first $k$ entries. No column in $Q-e$ has all of its first $k$ entries equal to zero, so we have $12^{k}-$ 1 such sets $F$.
For the sets $F$ where each column in $F$ has exactly one non-zero entry in its first $k$ entries, the root vectors of the columns in $F$ are the columns of the identity matrix of size $r-k$, so each such set $F$ has size $r-k$. For the remaining $2^{k}-(k+1)$ sets $F$, the root vectors of the columns in $F$ are the columns of the identity matrix of size $r-k$, along with the zero column, which has all of its entries equal to zero. The size of each of these sets $F$ is $r-k+1$. It follows that the size of $M$ is $2^{k}(r-k+1)$.
Note that there is no zero column in $P$. Moreover, since $r\geq 2$ and $k\leq r-2$, all of the columns in $P$ are distinct, so $M$ is simple. Observe that if we add fewer than $r-k$ columns to $e$, then we cannot obtain the zero column so all circuits containing $e$ in $M$ have size greater than $r-k$. Therefore $e$ is $k$-loose in $M$. ∎
3. $k$-loose elements in $GF(q)$-matroids
Zaslavsky and the author in [6] showed that a ternary matroid with no coloops that contains a $1$-loose element has linear size relative to its rank. In particular they proved the following.
Theorem 3.1.
For $r\geq 5$, let $M$ be a simple ternary matroid of rank $r$ containing a loose element and with no coloops. Then $|E(M)|\leq\lfloor\frac{41r-101}{2}\rfloor$ if $r>10$. If $r\leq 10$, then $|E(M)|\leq\lfloor\frac{35r-35}{2}\rfloor$.
Furthermore, they proved that if a cosimple ternary matroid contains two $1$-loose elements, then the rank of the matroid is bounded. This result was shown to hold for all $GF(q)$-matroids.
Theorem 3.2.
Let $M$ be a simple $GF(q)$-matroid with no coloops that has two $1$-loose elements $e$ and $f$. Then $r(M)\leq 2q$, or $\{e,f\}$ is a cocircuit of $M$.
We extend Theorem 3.2 to $k$-loose elements and prove Theorem 1.2. The following is the restatement of Theorem 1.2.
Theorem.
Let $M$ be a simple $GF(q)$-matroid with no coloops that has two $k$-loose elements $e$ and $f$. Then $r(M)\leq(q+1)(k-1)+2q$, or $\{e,f\}$ is a cocircuit of $M$.
Proof.
We may assume that $\{e,f\}$ is not a cocircuit of $M$. Therefore, there is a basis $B$ of $M$ that is disjoint from $\{e,f\}$. Note that, in the standard $GF(q)$-representation $P=[I_{r}|Q]$ of $M$ with respect to $B$, the column $e^{P}$ corresponding to $e$ and the column $f^{P}$ corresponding to $f$ have at most $k$ zeroes. By row and column scaling, we may assume that $e^{P}$ has all ones except that the first $k$ entries may be zero.
Let $f^{P^{\prime}}$ be the column obtained by removing the first $k$ entries of $f^{P}$. Suppose $f^{P^{\prime}}$ has $k+2$ non-zero entries that are all equal. Then $f^{P},e^{P}$, and at most $r-(k+2)$ columns from $I_{r}$ can be combined linearly over $GF(q)$ to obtain a column that has all its entries equal to zero. This would create a circuit of $M$ that contains $\{e,f\}$, and has size at most $r-k$, a contradiction. Therefore, $f^{P^{\prime}}$ has at most $k+1$ non-zero entries that are equal. Since the field $GF(q)$ has $q-1$ non-zero elements, it follows that $f^{P^{\prime}}$ has at most $(k+1)(q-1)$ non-zero entries. Additionally, $f^{P^{\prime}}$ has at most $k$ zeroes so it has at most $kq+q-1$ entries. Therefore, $f^{P}$ has at most $kq+q-1+k$ entries so the rank of $M$ is at most $(q+1)(k-1)+2q$. ∎
Corollary 3.3.
Let $M$ be a simple and cosimple $GF(q)$-matroid that has two $k$-loose elements. Then $r(M)\leq(q+1)(k-1)+2q$.
The following is a restatement of Corollary 1.3. It follows from Theorem 1.2 and provides the bound whose existence was established by Rajpal [5, Proposition 8].
Corollary.
Let $M$ be a simple $GF(q)$-matroid with no coloops that is $k$-paving. Then $M$ is a circuit, or $r(M)\leq(q+1)(k-1)+2q$.
Proof.
Note that every element of $M$ is $k$-loose. If $r(M)>(q+1)(k-1)+2q$, it follows from Theorem 1.2 that every pair of elements of $M$ is in a cocircuit of size two. It follows that $M$ is a circuit. ∎
The following special cases of Corollary 1.3 are noteworthy. The binary case was shown by Rajpal [5, Proposition 6].
Corollary 3.4.
Let $M$ be a simple binary matroid with no coloops that is $k$-paving. Then $M$ is a circuit, or $r(M)\leq 3k+1$.
Corollary 3.5.
Let $M$ be a simple ternary matroid with no coloops that is $k$-paving. Then $M$ is a circuit, or $r(M)\leq 4k+2$.
4. $k$-paving binary matroids
Theorem 1.1 bounds the size of a binary matroid with a $k$-loose element and rank at least $23k+$ 2. By Corollary 3.4, the rank of binary $k$-paving matroids is at most $13k+$ 1. Theorem 1.5 bounds the size of these matroids. The following is its restatement.
Theorem.
Let $M$ be a simple binary $k$-paving matroid of rank $r$, where $M$ is not a circuit and $r\geq k+4$, and let $t=\lfloor\frac{3k+1-r}{2}\rfloor$. Then $r(M)\leq 3k+1$, and
$|E(M)|\leq(r+1)+\sum_{i=0}^{t}\binom{k}{i}.$
Proof.
By Corollary 3.4, it is evident that $r(M)\leq 3k+1$. Suppose that $M$ has a circuit $C$ of size $r-k+1$. For an element $e$ of $C$, let $B$ be a basis of $M$ containing $C-e$. Consider the standard binary representation $P=[I_{r}|Q]$ of $M$ with respect to the basis $B$.
4.0.1.
Let $h$ be the sum of $m$ columns in $Q$. Then $h$ has at most $k+(m-1)$ zeroes. In particular, for any two columns $a$ and $b$ in $Q$, the sum $a+b$ has at most $k+1$ zeroes.
Assume that $h$ has at least $k+m$ entries that are zero. Then, at most $r-(k+m)$ columns in $I_{r}$ can be added to $h$ to obtain the column with all entries equal to zero. This would result in a circuit of size at most $r-k$ in $M$, a contradiction. Therefore 4.0.1 holds.
Note that the column $e^{P}$ in $Q$ corresponding to the element $e$ of $M$ has exactly $k$ zeroes. By reordering rows and columns of $P$, we may assume that the first $k$ entries of $e^{P}$ are zero. We denote the set of indices of the zero entries of a column $f^{P}$ in $Q$ by $\text{zeroes}(f^{P})$. The entries of $f^{P}$ excluding the first $k$ entries are the root entries of $f^{P}$. Let $g^{P}$ be a column in $Q-e^{P}$ such that $|\text{zeroes}(e^{P})\cap\text{zeroes}(g^{P})|=l$ is maximum.
4.0.2.
$r\leq 3k+1-2l.$
Observe that $g^{P}$ has at most $k$ entries that are zero; otherwise, $M$ has a circuit of size less than $r-k$. By 4.0.1, it follows that $g^{P}$ has at most $k-l$ root entries that are zero. Furthermore, by 4.0.1, at most $k+1-l$ root entries of $g^{P}$ are equal to one. Therefore, $g^{P}$ has at most $k+(k-l)+(k+1-l)=3k+1-2l$ entries. Thus, 4.0.2 holds.
Let $F$ be a set of all columns in $Q-e^{P}$ such that, for any pair $f^{P},g^{P}$ in $F$, the sum $f^{P}+g^{P}$ has its first $k$ entries equal to zero. Evidently, the first $k$ entries of all the columns in $F$ are identical.
4.0.3.
$|F|\leq 1.$
Suppose not, and let $a^{P}$ and $b^{P}$ be two columns in $F$. Note that the first $k$ entries of $a^{P}+b^{P}$ are all zero, so by 4.0.1, at most one root entry of $a^{P}+b^{P}$ is zero. It follows that $a^{P}+b^{P}+e^{P}$ has at most one non-zero entry. Therefore $M$ has a circuit of size at most four. Since $r\geq k+4$ and every circuit is of size greater than $r-k$, this is a contradiction. Thus, 4.0.3 holds.
4.0.4.
$|E(M)|\leq(r+1)+\sum_{i=0}^{l}\binom{k}{i}$.
Observe that each column in $Q-e^{P}$ has at most $l$ zeroes among the first $k$ entries. It follows that there are at most $\sum_{i=0}^{l}\binom{k}{i}$ sets $F$ in $Q-e^{P}$. Therefore, by 4.0.3, $|E(M)|\leq(r+1)+\sum_{i=0}^{l}\binom{k}{i}$ so 4.0.4 holds.
By 4.0.2, $l\leq\lfloor\frac{3k+1-r}{2}\rfloor=t$ so by 4.0.4, the result follows.
If $M$ has no circuit of size $r-k+1$, let $w$ be the smallest number such that $M$ has a circuit of size $r-w+1$. Clearly, $w<k$. By the same argument, it follows that $|E(M)|\leq(r+1)+\sum_{i=0}^{s}\binom{w}{i}$, where $s=\lfloor\frac{3w+1-r}{2}\rfloor$. Since $w<k$, it follows that $s\leq t$, and $|E(M)|\leq(r+1)+\sum_{i=0}^{t}\binom{k}{i}$. ∎
We apply Theorem 1.5 to $3$-paving binary matroids and prove Corollary 1.6.
Proof of Corollary 1.6.
It is clear that $r(M)\leq 10$. Note that if $r(M)$ is $7$, then $t$ equals $1$ so by Theorem 1.5, $|E(M)|$ is at most $12$. Similarly, if $r(M)=8,9,$ and $10$, then $|E(M)|$ is at most $13,11,$ and $12$, respectively. ∎
Acknowledgment
The author thanks James Oxley for helpful suggestions.